OpenStudy (anonymous):

I really need help understanding how to solve this Mathematics problem. s=125/1-0.7(y-1983) S=17 and y = 1992 Can someone Please Help Me.

OpenStudy (anonymous):

you need someone tho help you plug values into an equation??

OpenStudy (anonymous):

That and solving the equation.

OpenStudy (saifoo.khan):

\[s=\frac{125}{1-0.7(y-1983)}\]Like that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

are you sure your question is complete? or can u give the complete answer>

OpenStudy (anonymous):

the equation is like you show with this added y>=1983

OpenStudy (anonymous):

S will =17 and y is 1992

OpenStudy (saifoo.khan):

i still dont get it sorry. can u rephrase?

OpenStudy (anonymous):

Let me put it this way then. S=17 Y=1992 the equation is 17= 125/1+0.7(1992-1983), y.=1983 need to solve and not understanding.

OpenStudy (saifoo.khan):

Ohh!! so solving this, \[17 = \frac{125}{1+0.7(9)}\] \[17 = \frac{125}{1+6.3}\] \[17 = \frac{125}{7.3}\]

OpenStudy (saifoo.khan):

17 = 17.123

OpenStudy (anonymous):

Thank you. I was performing the figures in the equation in the wrong way. This will help me understand what I need to do better. Thank you, Thank you, Thank you.

OpenStudy (saifoo.khan):

No problem!! =) i can help u with other math stuff as well. if u are stuck anywhere!!

OpenStudy (anonymous):

Not at this moment. This is part of a team project for school and no one else was able to work the equation either. I will get back to the project now and let everyone else know how it works now that I know. If I get stuck again I will log back in and ask for the help. Again thank you.

OpenStudy (saifoo.khan):

Cool. Sure anytime.. you welcome..

OpenStudy (saifoo.khan):

PLease hit good answer. on the right.